Solve the nonlinear system of equations. Solve by using elimination or substitution methods.
x^+y^=5
x+y=3

Answer by Alan3354(30935)

What does x^ mean?
If it's x^2:
-------------
y = 3-x
Sub for y in the 1st eqn
x^2 + (3-x)^2 = 5
2x^2 - 6x + 9 = 5
x^2 - 3x + 2 = 0
(x-1)*(x-2) = 0
x = 1, y = 2 --> (1,2)
x = 2, y = 1 --> (2,1)
-------------
Use x^2 for x squared, x^3 for x cubed, etc